# Large Language Model-Assisted UAV Operations and Communications: A Multifaceted Survey and Tutorial

### v1

@@ Abstract
Uncrewed Aerial Vehicles (UAVs) are widely deployed across diverse applications due to their mobility and agility. Recent advances in Large Language Models (LLMs) offer a transformative opportunity to enhance UAV intelligence beyond conventional optimization-based and learning-based approaches. By integrating LLMs into UAV systems, advanced environmental understanding, swarm coordination, mobility optimization, and high-level task reasoning can be achieved, thereby allowing more adaptive and context-aware aerial operations. This survey systematically explores the intersection of LLMs and UAV technologies and proposes a unified framework that consolidates existing architectures, methodologies, and applications for UAVs. We first present a structured taxonomy of LLM adaptation techniques for UAVs, including pretraining, fine-tuning, Retrieval-Augmented Generation (RAG), and prompt engineering, along with key reasoning capabilities such as Chain-of-Thought (CoT) and In-Context Learning (ICL). We then examine LLM-assisted UAV communications and operations, covering navigation, mission planning, swarm control, safety, autonomy, and network management. After that, the survey further discusses Multimodal LLMs (MLLMs) for human–swarm interaction, perception-driven navigation, and collaborative control. Finally, we address ethical considerations, including bias, transparency, accountability, and Human-in-the-Loop (HITL) strategies, and outline future research directions. Overall, this work positions LLM-assisted UAVs as a foundation for intelligent and adaptive aerial systems.
@@ I Introduction
Uncrewed Aerial Vehicles (UAVs) have become a salient enabling technology across a wide range of applications, including environmental monitoring [74], public safety [105], transportation [85], agriculture [92], and infrastructure inspection [4]. For example, UAVs can track environmental changes, support emergency response during disasters, assess crop health, deliver parcels, and assist with traffic management. Due to their flexibility and mobility, UAVs provide efficient and cost-effective solutions in situations where conventional ground-based systems are limited or infeasible [56].
From a communication perspective, UAVs are increasingly important for enhancing wireless networks. They can serve as flying base stations to expand network coverage, especially in areas with sparse or damaged infrastructure [40, 3]. UAVs can also operate as relays [64], [95], collect data from ground devices [63], and provide edge computing capabilities in the air [65], reducing communication latency. On the other hand, UAVs may operate as cellular users, requiring reliable, low-latency connectivity for safe navigation, control, and sensing [6]. When multiple UAVs cooperate, they can form Flying Ad hoc Networks (FANETs) that self-organize and provide either extended range [94] or communication services over large geographical areas or in environments without pre-existing infrastructure. These capabilities make UAVs a critical component of wireless backhauling, smart city applications, and future mobile communication systems [87].
@@ I-A Why Combining UAVs and LLMs Matters
Compared with conventional ML-based optimization methods, LLMs introduce a fundamentally different paradigm centered on semantic reasoning rather than parameter tuning alone. While traditional ML and DRL approaches optimize numerical objectives through repeated interaction with simulated environments, LLMs operate at a higher abstraction level, enabling structured reasoning over heterogeneous inputs and mission-level objectives. This capability allows LLMs to function not merely as optimizers, but as cognitive coordinators within UAV systems. Specifically, in UAV communications and networking scenarios, LLMs can interpret network logs, topology information, channel conditions, and mission constraints in natural language form, transforming raw telemetry into structured situational awareness. Instead of solving isolated optimization subproblems (e.g., power allocation or trajectory refinement), LLMs can synthesize cross-layer information to guide adaptive spectrum management, cooperative relay strategies, and swarm-level resource coordination. By reasoning over operational rules, safety constraints, and mission priorities, LLMs can generate context-aware decision strategies that complement lower-level control modules.
Beyond communication optimization, LLMs enhance UAV autonomy by bridging perception, planning, and execution. Through multimodal integration, LLMs can align sensor data, map representations, and textual mission descriptions into a unified reasoning framework. This enables dynamic task decomposition, contingency planning, and collaborative swarm coordination under uncertain or partially observable conditions. Rather than relying solely on fixed reward functions, LLM-driven systems can adjust strategies based on evolving mission semantics and operator intent. In addition, this reduces the dependence on handcrafted rule sets and specialized pipelines and mitigates the need for multiple task-specific models. By unifying diverse data modalities—natural-language commands, telemetry streams, network metrics, and environmental semantics—LLMs provide a coherent operational layer that supports complex, multi-objective UAV missions.
@@
In the following, we present a unified abstraction at the system-level to explain how LLMs interact holistically with UAV communication, control, and decision-making. This is a macro-level methodological framework for understanding LLM–UAV integration that is organized into three layers:
@@
  • The control and communication layer, at the bottom, translate high-level plans and semantic decisions into executable actions, such as control commands, flight trajectories, and communication policies. Execution outcomes and system states are abstracted and relayed to higher layers, forming a closed-loop learning and adaptation process.
@@ I-B Main Contributions
This paper presents a comprehensive investigation of LLM-UAV interactions. It systematically examines how LLM-assisted techniques, such as Retrieval-Augmented Generation (RAG) and prompt engineering, are fundamentally transforming UAV operations by enhancing decision-making, enabling adaptive path planning, and supporting real-time human–UAV interaction. In addition, we explore the paradigm shift from text-only LLMs to MLLMs, analyzing their architectures, training methodologies, and emerging applications in vision–language navigation, swarm coordination, and high-level mission planning. After that, beyond technical advancements, we critically assess the ethical implications of integrating LLMs into UAV operation and communications, including privacy and data security, bias and fairness, transparency and accountability, and environmental sustainability. Finally, we present an LLM-aided UAV case study and identify the future research directions. Overall, this work serves as a comprehensive technical survey and tutorial, as well as an essential ethical guide for the evolving landscape of LLM-assisted autonomous UAV systems.
@@
  • In-depth discussion of MLLMs as a transformative enabler for next-generation UAV systems, detailing their core architectures, training paradigms, and emerging capabilities such as Multimodal Chain of Thought (MCoT) reasoning and Multimodal ICL (M-ICL). We further demonstrate how MLLMs enable vision–language navigation and facilitate coordinated swarm control, underscoring their growing importance in complex perception, reasoning, and decision-making tasks for autonomous UAVs.
  • Rigorous analysis of the ethical challenges associated with deploying LLM-assisted UAV systems (LAUS), including risks related to bias and fairness, transparency and accountability in autonomous decision-making, and the environmental impact of LLMs. Beyond identifying these issues, the survey overviews related mitigation strategies, emphasizing LLM observability for systematic monitoring and debugging. Together, these recommendations establish a practical foundation for building trustworthiness, responsibility, and resilience in LLM-assisted UAV communications.
@@ I-C Survey of Surveys
Table I summarizes key surveys on integrating LLMs with UAVs. Early comprehensive research by Javaid et al. [49] examines how LLMs can enhance UAV autonomy, intelligence, and operational efficiency. Their survey reviews a wide range of notable LLM architectures, including Bidirectional Encoder Representations from Transformers (BERT) [21], Generative Pretrained Transformers (GPT), T5, XLNet, and evaluates their applicability to various UAV tasks such as natural language-based control, real-time data analysis, and mission planning. This work lays the foundation for understanding LLM-aided UAV’s technical potential.
@@
[16] Analyzes 17 papers to map the research landscape (2021–2024). Applications: UAV Optimization (35%), Adaptation (40%), Communication (25%). Models: GPT most used (58%). Highlights challenges including hardware limitations, latency, security, and ethical alignment.
@@
This survey LLMs and MLLMs that enhance UAV autonomy in navigation, communication, and swarm coordination, using techniques such as fine-tuning, RAG, and prompt engineering. The survey addresses ethical issues, including bias and transparency, and recommends HITL oversight to ensure safety. Practical and future research directions are provided. It positions LLM-assisted UAVs as a transformative, intelligent aerial framework. LLMs enhance UAV navigation, path planning, and real-time swarm coordination, surpassing traditional optimization methods. In communications, they optimize spectrum management, resource allocation, and the integration of 5G/6G networks. MLLMs enable multimodal tasks like search and rescue by fusing visual, sensor, and language data. The technology also addresses ethical deployment through safety verification and human oversight frameworks.
@@
Unlike existing surveys, our work offers a focused and in-depth examination of UAV communications, emphasizing how LLMs can enhance communication reliability, autonomy, and decision-making in this critical field. We systematically analyze the LLM functionalities most relevant to UAV communication systems and provide detailed insights into their practical adoption. We also explore the emerging role of MLLMs in integrating visual, sensor, and textual data to enable advanced communication-aware UAV applications. Beyond technical aspects, we address ethical and societal implications, including safety, privacy, accountability, and potential misuse of LAUS. To bridge theory and practice, we present a detailed case study demonstrating how LLM-assisted ICL can support real-world UAV communication scenarios. Finally, we discuss emerging trends and open research directions, positioning our survey as a timely and distinctive contribution to the evolving field of intelligent UAV communications.
@@ I-D Organization of This Survey
The remainder of this paper is systematically structured as follows. Section II details the methodologies for LLM training, adaptation (including fine-tuning, RAG, and prompt engineering), and deployment strategies tailored for UAV operations. Section III explores the application of LLMs specifically in UAV communications, covering navigation, swarm control, and network optimization. Section IV examines the transformative role of MLLMs, detailing their architectures and their application in Vision-Language Navigation (VLN) and human-swarm interaction. Section V critically examines the associated ethical, safety, and societal implications and associated mitigation frameworks. Section VI presents a case study on LLM-assisted ICL for multi-UAV data collection and velocity control. Section VII outlines emerging trends and future research directions, and finally Section VIII concludes the survey. Fig. 1 provides the general organization of the survey.
@@ II LLM Training, Adaptation, and Deployment for UAV Operations
II-A LLM Model Architectures
II-A1 Encoder-Only Architectures
These are LLM architectures designed to generate rich contextual representations of input data. Their core mechanism is bidirectional self-attention, which enables them to capture a word’s context by considering all surrounding words. This capability makes them particularly effective for tasks that require deep language comprehension, such as understanding objects and scenes. The development of this architecture was pioneered by BERT, which introduced the masked language modeling objective.
In UAV-related applications, encoder-only models have demonstrated strong performance in text understanding tasks. In this line, Silalahi et al. [106] propose a Transformer-based Named Entity Recognition (NER) framework for extracting forensic information from UAV flight logs. Using a manually annotated DJI dataset with six entity types, their BERT-based architecture achieved an F1 score of 91.348%, enabling efficient identification of forensically relevant entities.
II-A2 Encoder-decoder Architectures
These employ a three-module design to handle complex tasks: an encoder that processes the input, a decoder that generates the output, and a cross-attention mechanism that enables the decoder to focus on relevant parts of the input. This structured division of responsibilities often improves the sample and inference efficiency, making it well-suited for models such as T5, BART [58], and OpenBA [132].
The advanced text understanding and generation capabilities of the T5 language model provide a powerful means to enhance the autonomy and utility of UAVs significantly. By enabling natural language command interpretation and automated data summarization, T5 can streamline human-UAV interaction and convert raw sensor data into actionable intelligence, increasing efficiency in critical missions such as disaster response [49]. Encoder-decoder LLMs offer advantages in data efficiency, especially when tasks require both understanding and generative capabilities [71].
II-A3 Decoder-Only Architectures
These are a foundational class of LLM architectures designed for generative tasks, producing human-like text that is coherent and contextually relevant [96]. Their core mechanism is causal (autoregressive) self-attention, which restricts the model to attending only to previous tokens in a sequence when generating the next word. This enables step-by-step text generation while maintaining contextual consistency. The modern era of decoder-only models began with the GPT series, which established the paradigm of pretraining on large-scale text corpora to achieve advanced language-generation capabilities. Today, decoder-only architectures underpin many influential Generative AI (GAI) systems, from research-focused models to leading commercial applications.
In UAV-related applications, decoder-only LLMs have shown substantial potential for natural language interaction and control. For example, Tazir et al. [109] integrated OpenAI’s ChatGPT with the PX4/Gazebo UAV simulation platform, enabling intuitive UAV control through spoken or written commands. Their framework relies on prompt-engineered command generation with a verification mechanism to ensure safety and reliability, achieving high command-generation accuracy in zero-shot settings and demonstrating the feasibility of user-friendly human–UAV interfaces. Additionally, Zhang et al. [134] introduce Open Pretrained Transformers (OPT), a suite of decoder-only models with up to 175 billion parameters. OPT-175B achieves performance comparable to GPT-3 while training with a significantly lower carbon footprint, thereby supporting reproducible research. In general, these studies highlight the versatility of decoder-only LLMs for natural-language-driven UAV control and automation, while also underscoring their reliance on large-scale training data.
Overall, the architecture of an LLM, whether encoder-only, encoder-decoder, or decoder-only, directly affects how effectively it can leverage pre-training to acquire language understanding, reasoning, and generative capabilities. Pretraining converts these architectural potentials into practical intelligence by exposing the model to large-scale general and domain-specific datasets, enabling it to develop emergent abilities essential for UAV applications. The following subsection describes the pretraining process, data preparation strategies, and the challenges of adapting LLMs for UAV-specific contexts.
II-B LLM Pre-training
The pretraining of LLMs primarily relies on next-token prediction, which gives models strong language comprehension, reasoning, and long-range dependency modeling capabilities. In UAV communications, pretraining typically uses two types of data, namely general corpora and UAV-specific datasets. General corpora (e.g., Wikipedia, technical blogs, and web content) provide linguistic diversity and support understanding of technical documentation, whereas specialized datasets focus on UAV- and aerial-specific knowledge. These include $3^{rd}$ Generation Partnership Project (3GPP) standards, Request For Comments (RFC) documents, academic literature, network configuration logs, protocol specifications, simulation reports, and operational UAV repositories. Incorporating such domain-specific sources enhances the model’s understanding of UAV systems, communication architectures, and optimization tasks, making it better suited to UAV-oriented reasoning.
Pretraining LLMs from scratch for UAV applications can be highly resource-intensive. UAV-relevant data, such as flight logs, sensor readings, communication traces, and environmental information, require careful handling in domain-specific formats, including navigation equations, flight dynamics, and swarm coordination rules. Moreover, the computational cost of large-scale training poses significant challenges for real-time UAV systems and edge deployments. For instance, training a frontier-scale model such as LLaMA 3.1 (405B) requires tens of millions of GPU hours and incurs costs on the order of tens of millions of dollars [23]. Consequently, complete pretraining is often infeasible in UAV networks, underscoring the importance of fine-tuning pretrained models to achieve cost-effective and practical domain adaptation.
II-C LLM Fine-tuning

### v2

# arXiv:2602.19534v2[cs.RO] 01 Mar 2026

@@ Abstract
Uncrewed Aerial Vehicles (UAVs) are widely deployed across diverse applications due to their mobility and agility. Recent advances in Large Language Models (LLMs) offer a transformative opportunity to enhance UAV intelligence beyond conventional optimization-based and learning-based approaches. By integrating LLMs into UAV systems, advanced environmental understanding, swarm coordination, mobility optimization, and high-level task reasoning can be achieved, thereby allowing more adaptive and context-aware aerial operations. This survey systematically explores the intersection of LLMs and UAV technologies and proposes a unified framework that consolidates existing architectures, methodologies, and applications for UAVs. We first present a structured taxonomy of LLM adaptation techniques for UAVs, including pretraining, fine-tuning, Retrieval-Augmented Generation (RAG), and prompt engineering, as well as key reasoning capabilities such as Chain-of-Thought (CoT) and In-Context Learning (ICL). We then examine LLM-assisted UAV communications and operations, covering navigation, mission planning, swarm control, safety, autonomy, and network management. Subsequently, the survey discusses Multimodal LLMs (MLLMs) for human–swarm interaction, perception-driven navigation, and collaborative control. Finally, we discuss ethical considerations, including bias, transparency, accountability, and Human-in-the-Loop (HITL) strategies, and outline directions for future research. Overall, this work positions LLM-assisted UAVs as a foundation for intelligent and adaptive aerial systems.
@@ I Introduction
Uncrewed Aerial Vehicles (UAVs) have become a salient enabling technology across a wide range of applications, including environmental monitoring [66], public safety [91], transportation [75], agriculture [79], and infrastructure inspection [6]. For example, UAVs can track environmental changes, support emergency response during disasters, assess crop health, deliver parcels, and assist with traffic management. Due to their flexibility and mobility, UAVs provide efficient and cost-effective solutions in situations where conventional ground-based systems are limited or infeasible [52, 4, 3].
From a communication perspective, UAVs are increasingly important for enhancing wireless networks. They can serve as flying base stations to expand network coverage, especially in areas with sparse or damaged infrastructure [37, 5]. UAVs can also operate as relays [59], [82], collect data from ground devices [58], and provide edge-computing capabilities in the air [60], thereby reducing communication latency. On the other hand, UAVs may operate as cellular users, requiring reliable, low-latency connectivity for safe navigation, control, and sensing [8]. When multiple UAVs cooperate, they can form Flying Ad hoc Networks (FANETs) that self-organize and provide either extended range [81] or communication services over large geographical areas or in environments without pre-existing infrastructure. These capabilities make UAVs a critical component of wireless backhauling, smart city applications, and future mobile communication systems [76].
@@ I-A Why Combining UAVs and LLMs Matters
Compared with conventional ML-based optimization methods, LLMs introduce a fundamentally different paradigm centered on semantic reasoning rather than parameter tuning alone. While traditional ML and DRL approaches optimize numerical objectives through repeated interactions with simulated environments, LLMs operate at a higher level of abstraction, enabling structured reasoning over heterogeneous inputs and mission-level objectives. This capability allows LLMs to function not merely as optimizers but as cognitive coordinators within UAV systems. Specifically, in UAV communications and networking scenarios, LLMs can interpret network logs, topology information, channel conditions, and mission constraints in natural-language form, thereby transforming raw telemetry into structured situational awareness. Instead of solving isolated optimization subproblems (e.g., power allocation or trajectory refinement), LLMs can synthesize cross-layer information to guide adaptive spectrum management, cooperative relay strategies, and swarm-level resource coordination. By reasoning over operational rules, safety constraints, and mission priorities, LLMs can generate context-aware decision strategies that complement lower-level control modules.
Beyond communication optimization, LLMs enhance UAV autonomy by bridging perception, planning, and execution. Through multimodal integration, LLMs can align sensor data, map representations, and textual mission descriptions into a unified reasoning framework. This enables dynamic task decomposition, contingency planning, and collaborative swarm coordination under uncertain or partially observable conditions. Rather than relying solely on fixed reward functions, LLM-driven systems can adapt strategies to evolving mission semantics and operator intent. In addition, this reduces the dependence on handcrafted rule sets and specialized pipelines and mitigates the need for multiple task-specific models. By unifying diverse data modalities—natural-language commands, telemetry streams, network metrics, and environmental semantics—LLMs provide a coherent operational layer that supports complex, multi-objective UAV missions.
@@
In the following, we present a unified system-level abstraction that explains how LLMs interact holistically with UAV communication, control, and decision-making. This is a macro-level methodological framework for understanding LLM–UAV integration that is organized into three layers:
@@
  • The control and communication layer, at the bottom, translates high-level plans and semantic decisions into executable actions, such as control commands, flight trajectories, and communication policies. Execution outcomes and system states are abstracted and relayed to higher layers, forming a closed-loop learning and adaptation process.
@@ I-B Main Contributions
This paper presents a comprehensive investigation of LLM-UAV interactions. It systematically examines how LLM-assisted techniques, such as Retrieval-Augmented Generation (RAG) and prompt engineering, are fundamentally transforming UAV operations by enhancing decision-making, enabling adaptive path planning, and supporting real-time human–UAV interaction. In addition, we explore the paradigm shift from text-only LLMs to MLLMs, analyzing their architectures, training methodologies, and emerging applications in vision–language navigation, swarm coordination, and high-level mission planning. Beyond technical advancements, we critically assess the ethical implications of integrating LLMs into UAV operations and communications, and identify directions for future research. Overall, this work serves as a comprehensive technical survey and tutorial, as well as an essential ethical guide for the evolving landscape of LLM-assisted autonomous UAV systems.
@@
  • In-depth discussion of MLLMs as a transformative enabler for next-generation UAV systems, detailing their core architectures, training paradigms, and emerging capabilities such as Multimodal Chain of Thought (MCoT) reasoning and Multimodal ICL (M-ICL). We further demonstrate that MLLMs enable vision–language navigation and facilitate coordinated swarm control, underscoring their growing importance for complex perception, reasoning, and decision-making tasks in autonomous UAVs.
  • Rigorous analysis of the ethical challenges associated with deploying LLM-assisted UAV systems (LAUS), including risks related to bias and fairness, transparency and accountability in autonomous decision-making, and the environmental impact of LLMs. Together, these analyses establish a practical foundation for building trustworthiness, responsibility, and resilience in LLM-assisted UAV communications.
@@ I-C Survey of Surveys
Table I summarizes key surveys on integrating LLMs with UAVs. Early comprehensive research by Javaid et al. [46] examines how LLMs can enhance UAV autonomy, intelligence, and operational efficiency. Their survey reviews a wide range of notable LLM architectures, including Bidirectional Encoder Representations from Transformers (BERT) [22], Generative Pretrained Transformers (GPT), T5, XLNet, and evaluates their applicability to various UAV tasks such as natural language-based control, real-time data analysis, and mission planning. This work lays the foundation for understanding the technical potential of LLM-aided UAVs.
@@
[17] Analyzes 17 papers to map the research landscape (2021–2024). Applications: UAV Optimization (35%), Adaptation (40%), Communication (25%). Models: GPT most used (58%). Highlights challenges, including hardware limitations, latency, security, and ethical alignment.
@@
This survey LLMs and MLLMs that enhance UAV autonomy in navigation, communication, and swarm coordination, using techniques such as fine-tuning, RAG, and prompt engineering. The survey discusses ethical issues such as bias and transparency. Practical and future research directions are provided. It positions LLM-assisted UAVs as a transformative, intelligent aerial framework. LLMs enhance UAV navigation, path planning, and real-time swarm coordination, surpassing traditional optimization methods. In communications, they optimize spectrum management, resource allocation, and the integration of 5G/6G networks. MLLMs enable multimodal tasks like search and rescue by fusing visual, sensor, and language data. The technology also addresses ethical deployment through safety verification and human oversight frameworks.
@@
Unlike existing surveys, our work provides a focused, in-depth examination of UAV communications, emphasizing how LLMs can enhance communication reliability, autonomy, and decision-making in this critical field. We systematically analyze the LLM functionalities most relevant to UAV communication systems and provide detailed insights into their practical adoption. We also explore the emerging role of MLLMs in integrating visual, sensor, and textual data to enable advanced communication-aware UAV applications. Beyond technical aspects, we analyze ethical and societal implications, including safety, privacy, accountability, and potential misuse of LAUS, and discuss emerging trends and open research directions, positioning our survey as a timely and distinctive contribution to the evolving field of intelligent UAV communications.
@@ I-D Organization of This Survey
The remainder of this paper is systematically structured as follows. Fig. 1 provides the general organization of the survey. Section II details the methodologies for LLM pre-training, adaptation (including fine-tuning, RAG, and prompt engineering), and deployment strategies tailored for UAV operations. Section III explores the application of LLMs specifically in UAV communications, covering navigation, swarm control, and network optimization. Section IV examines the transformative role of MLLMs, detailing their architectures and their application in Vision-Language Navigation (VLN) and human-swarm interaction. Section V critically examines the associated ethical implications and future research directions, and finally Section VI concludes the survey.
@@ II LLM Training, Adaptation, and Deployment for UAV Operations
II-A LLM Model Architectures and Pre-training
Large language models can be broadly categorized into three architectural paradigms: encoder-only, encoder–decoder, and decoder-only. Encoder-only models (e.g., BERT) rely on bidirectional self-attention to produce deep contextual representations, making them particularly effective for language understanding and classification tasks. Encoder–decoder models (e.g., T5, BART) separate input encoding and output generation with cross-attention, enabling efficient handling of tasks that require both comprehension and structured generation. Decoder-only models (e.g., GPT-style architectures) adopt causal self-attention for autoregressive text generation, excelling in open-ended reasoning and interactive language production. Their differences become clear in UAV applications. For example, in UAV log analysis, a BERT-based encoder-only model has been used for named entity recognition (NER) to extract forensic entities from DJI flight logs, achieving high F1 performance. This reflects its strength in structured understanding. In contrast, encoder–decoder models such as T5 are better suited for tasks like translating natural-language commands into structured UAV instructions or summarizing sensor data. Decoder-only models further extend this capability by enabling interactive UAV control, in which systems such as ChatGPT are integrated with PX4/Gazebo to generate executable flight commands via prompt-based reasoning.
Overall, the architectural choice determines whether the model primarily excels at understanding, understanding plus structured generation, or open-ended generative interaction, which directly shapes its role in UAV intelligence systems.
Pretraining equips LLMs with general language understanding and reasoning, primarily through next-token prediction on large-scale text corpora. This process enables models to capture long-range dependencies and acquire broad world knowledge before being adapted to specific domains. In UAV communications, pretraining typically combines general corpora (e.g., Wikipedia and technical web content) with UAV-specific sources such as 3GPP standards, RFC documents, academic papers, protocol specifications, and network logs. Incorporating these domain-relevant materials improves the model’s understanding of UAV architectures, communication protocols, and optimization tasks, thereby supporting more reliable UAV-oriented reasoning. However, training a large model from scratch for UAV applications is extremely resource-intensive. UAV data (e.g., flight logs, communication traces, and sensor records) are highly specialized and computationally expensive to process at scale. For example, training frontier-scale models such as LLaMA 3.1 (405B) requires massive GPU resources and prohibitive cost. As a result, full pretraining is typically impractical for UAV systems, making fine-tuning of pretrained models the more feasible and cost-effective approach for domain adaptation.
II-B LLM Fine-tuning for UAV Applications